OpenStudy (anonymous):

Describe how to transform the quantity of the third root of x to the fourth power, to the fifth power into an expression with a rational exponent

OpenStudy (inkyvoyd):

I could probably go a few more steps, but I believe you can take it from here :)

OpenStudy (anonymous):

Wouldn't it be x radical 20/3 ? Could it be simplified more?

OpenStudy (inkyvoyd):

I got the fifth power of x actually... where are you getting the radicals from?

OpenStudy (anonymous):

\[\sqrt[3]{x ^{4}}^{5}\]

OpenStudy (inkyvoyd):

Sigh.

OpenStudy (inkyvoyd):

Yup, you're right. :) think I read the problem wrong, and proceeded to type a huge jumbo mess in Tex >.<

OpenStudy (inkyvoyd):

though to describe the process you'd just have to do it step by step

OpenStudy (anonymous):

Okay. Could you help with one more?

OpenStudy (inkyvoyd):

sure, I guess

OpenStudy (anonymous):

I would do the same steps to simplify \[\left( \sqrt[3]{x}\right)^12\]

OpenStudy (inkyvoyd):

is that a one half or a 12?

OpenStudy (anonymous):

12 I couldn't get the 2 small.

OpenStudy (inkyvoyd):

okay, well yeah, you would just rewrite as \(\Huge (x^\frac{1}{3})^{12}=x^{\frac{1}{3}*12}\)

OpenStudy (inkyvoyd):

also, you may know this already, but a common mistake is to think that \(\Huge x^{(a^b)}=(x^a)^b\). The two are entirely different statements, though.

OpenStudy (anonymous):

Im so confused now. :(

OpenStudy (inkyvoyd):

starting where?

OpenStudy (anonymous):

Would it be x radical 4 then?

OpenStudy (inkyvoyd):

x to the power of 4 actually

OpenStudy (anonymous):

So what does radical mean?

OpenStudy (inkyvoyd):

a radical is an nth root. it's basically a square root sign you see, but instead with cube root (denoted with a small three), or a fourth root, or in general any nth root where n is a positive integer.

OpenStudy (anonymous):

Oh okay! Well thanks for all the help! Really appreciate it!

OpenStudy (inkyvoyd):

no problem - let me know if you need any more help!